OpenStudy (anonymous):

how do you solve a+5=-5a+5

OpenStudy (anonymous):

get the like terms on each of the sides. so get "a" terms on one side, and the constants on the other sides. a+5a=5-5

OpenStudy (anonymous):

a +5 = -5a + 5 - 5 - 5 a = -5a 6a = 0 a = 0 :)